2X x 20 5 pt. a. Find the cdf and pdf of Y in terms of the cdf and pdf of X. of Y when X is a Gaussian random variable with zero mean and variance-4
Answer #1

a. CDF of Y is FY(y) = P(Y-y) = P(3X-y) = Fx(y/3) if y < 0 = Fx(0) + Fx(y/2)-Fx(0) = Fx(y/2) if y > 0 PDF of Y is fy (y) = dF